OpenStudy (anonymous):

\[2x ^{1/2}-x ^{1/4}=3\] HELP

OpenStudy (anonymous):

This is a very simple case once you know how to do it. \[2x^.5=2*\sqrt{x}\] \[x ^{1/4}=\sqrt[4]{x}\] So if you enter this into a TI-89 using the solver you get: \[x=81/16\]